# SWaT: Statistical Modeling of Video Watch Time through User Behavior Analysis

### v1

@@ 3.3.2. Simplification: A Stationary Watching Probability
One may simplify the assumption on non-stationary watching probabilities in Section 3.3.1 by assuming a uniform watching probability over the entire video horizon, i.e., the user’s interest in continuing watching the video does not vary with the played time. As will be revealed later on and in Section 3.4, this simplified model is closely related to the weighted logistic regression (WLR, Covington et al., 2016).
@@
While Eq. 9 takes the same functional form as the watch-time estimator in WLR (Covington et al., 2016), it operates without extra assumptions or approximation in our framework. Concretely, Covington et al. (2016) also estimate watch time by the exponentiated logits (Eq. 9), but require the extra assumption of small click probability to establish certain approximation. Click probability is inapproprate to our target of short video platforms, as discussed in Sec. 1 and 2.1.
@@ 3.4. Connection with WLR
As a final remark, the functional similarity between ${\mathcal{L}}_{\mathrm{WLR}}$ and ${\mathcal{L}}_{\mathrm{v-geo}}$ indicates that WLR approximately models video watch time by a geometric distribution with uniform watching probability over the entire video horizon, which accounts for neither real user’s varying interests in continuing watching nor different video lengths. By contrast, the bucketization strategy employed in our framework can naturally cope with both desiderata should data support. For example, to account for the differences in video length, the model may learn unequal $\{\widehat{p}_{i}\}$ whose values become negligible when the index $i$ is greater than some smaller integer for short videos (due to their smaller lengths) and larger integer for long videos.
@@ 4.1. Public Dataset
CIKM16 Cup dataset contains sequences of anonymous transactions provide by DIGINETICA, and the goal is to predict the dwell time in search results for each session. The dataset contains $310,302$ sessions, a total of $122,991$ items and the average dwell time of each session is $3.981$. For each session, each item in this session is used as a feature in the model input.
@@
We compare our method with state-of-the-art methods for video watch time prediction, including Weighted Logistic Regression (WLR, Covington et al., 2016), Duration-Deconfounded Quantile Regression (D2Q, Zhan et al., 2022), Ordinal Regression (OR, Crammer and Singer, 2001), and Tree based Progressive Regression Model (TPM, Lin et al., 2023). A detailed description of each baseline method is provided in Appendix B.
@@
where $\mathrm{round}(\cdot)$ denotes rounding to nearest integer, $c=50$ for the KuaiRec dataset, and $c=100$ for the CIKM dataset, both of which are conveniently chosen based on the scale of the data target values. In inference time, the effect of multiplication by $c$ is reverted before calculating the evaluation metrics. On both datasets, we conveniently use $100$ buckets, defined by the $1$-percentiles of the data target values. Sec. 4.1.6 conducts ablation study on number of buckets, showing that our method is generally robust to this design parameter. We use AdamW optimizer (Loshchilov and Hutter, 2017) with weight decay 00.
@@
  • MAE: This metric measures absolute numerical error. Denote the true observed target value as $t_{i}$ and the predicted value as $\widehat{t}_{i}$, we have
@@
    where $N$ denotes test set size. Smaller MAE suggests better model performance.
@@
Table 1. Performance of our models on the metrics XAUC and MAE on KuaiRec and CIKM16 datasets. Baseline results are directed cited from Lin et al. (2023). The best result in each metric is bold, and the second best result is italic.
@@
We compare the results of our models and the baselines in Table 1, where the baseline results are directly cited from Lin et al. (2023). In Table 1, our two models SWaT-Binom and SWaT-Geo are consistently ranked top-2 across all the metrics and datasets. The better performance of our models over the SoTA method TPM indicates the efficacy of our user-centric perspective and statistical framework for modeling watch time. The gain of our methods over WLR may further testify our consideration on training stability and regularization, which are naturally incorporated in our proposed framework, as discussed in Section 3.4.
@@
From Fig. 4 and Fig. 5 and recall the best baseline results, it is clear that our models are generally robust to the number of buckets. While our convenient choice of $100$ buckets in Section 4.1.3 works well across the combination of models, metrics, and datasets, performance on a specific setting may be further improved by tuning the bucket numbers. For example, on the KuaiRec dataset, $50$ buckets suits SWaT-Geo model best, and using $20$ or even $10$ buckets can improve SWaT-Binom model’s XAUC. Using more buckets, such as $200$, however, generally does not provide much gain over using $100$ or fewer buckets. Hence, for a practical rule-of-thumb, one may start with $100$ buckets, and gradually reduce the bucket numbers while closely monitoring the desired metrics.
(b): Are the training of our models stable and well converged?
@@ 4.2. Industrial Offline Experiment
Due to the large size of the dataset, we train the models with one pass of the training set, i.e., for one epoch. Our models and the baseline use the same model architecture, except for the inevitable changes in the loss functions and the corresponding model output layer. All models use the same set of features and hyperparameters, both of which are optimized for the baseline model, which leaves room for further improving our models’ performance. Bucket definition is investigated in Section 4.2.6.
@@
As per industrial common practice, we evaluate the prediction of video watch time by the Pearson correlation between the predicted watch time and the ground truth. Large Pearson correlation indicates that higher predicted values mostly correspond to higher ground-truth values, and thus the prediction model can be more helpful for selecting videos to recommend.
@@ 4.2.5. Main Results
Table 2 compares the performance of our three models discussed in Section 3 against the baseline model. It is clear that all our proposals, even the un-bucketized vanilla geometric model, perform significantly better than the baseline, demonstrating the efficacy of our user-centric perspective and statistical framework for modeling watch time. As a side note, the simplest vanilla geometric model is less performant on other industrial topline metrics, which we omit details due to non-disclosure policy, and thus we will continue focusing on our two bucketized models.
@@
We investigate the performance of our SWaT-Binom model and SWaT-Geo model under various bucket definition. Our intention is to find a bucket definition that performs well while having fewer buckets. We hypothesize that due to the long-tail distribution of video watch time and our desiderata of modeling the right tail well, a dense uniform grid may not be desirable. With this motivation, we compare the following six choices of the bucket endpoints.
@@
Table 3. Gain over the baseline model of the SWaT-Binom and SWaT-Geo model under various choices of bucket endpoints listed in Section 4.2.6, together with the number of buckets under each choice. The gain is calculated as a direct subtraction by the baseline model’s result in Table 2. Highest numeric in each column is bold.
@@
Table 3 shows the gain of the SWaT-Binom model and SWaT-Geo model over the baseline model under these six definitions of bucket endpoints. We also tabulate the number of buckets under these six definitions. Since the bucket endpoints need to be distinct, we remove the repeated percentile values, and hence the number of buckets can be fewer than the arithmetically calculated counts. For example, in Choice 1, using the 5-percentiles to define the bucket endpoints should ideally give 20 buckets, but removing the repeated percentiles eventually gives us only 18 buckets.
@@ 4.3. Online Experiment
We conduct online A/B test by implementing our method onto the recommendation system on an real-world short video platform with hundreds of millions of daily-active users. Specifically, our experiment focuses on the performance of our method on long videos, which are more relevant to the event of video watch time.
@@
Table 4. Gain of our model over the production model on long videos in an online A/B test over five days. Reported metrics are accumulated watch time and accumulated watch count. For both metrics, the higher the results, the better. The gain is calculated as a direct subtraction by the baseline model’s result.
@@
Table 4 shows the results of online A/B test in each day over five days of traffic, on the metrics accumulated video watch time and accumulated watch counts, both of which are positive metrics, i.e., the higher the better. It is clear that our model provides significant gain on these two metrics in each day of the testing window. This result coincides with our industrial offline experiments (Section 4.2) and public datasets’ experiments (Section 4.1) that our method’s predicted video watch time is more aligned with the ground-truth watch time. Hence, our model can more accurately distinguish videos that attract the users most, and thus the recommendations based on our model increase users’ watch time and the number of watched videos (watch count).
@@ 5. Conclusion
Video watch time prediction is one of the central problems in (short) video recommendation, as it directly affects the recommendation quality and user experience. Unlike conventional binary prediction tasks, this problem is significantly challenged by the complex user-video interaction; while most of the existing literature focus on black-box mechanical enhancement of the classical regression/classification losses. In this paper, we for the first time take on a user-centric perspective to model video watch time, from which we propose a white-box statistical framework that translates our domain knowledge on users’ video-watching behavior into principled statistical models. To better model users’ fluctuating interests in continuing watching over the video horizon, we employ the bucketization strategy, which additionally helps to stabilize training objectives. Extensive experiments of various nature validate our perspective and derived models, showing that our predictions align well with the ground-truth watch time.
@@ Appendix A The Expectation of SWaT-Geo Model
Recall from Section 3.3.1 that for a given set of continuation probabilities $\left\{p_{i}\right\}_{i=1}^{N+1}$, if the observation $t\in B_{i}=(x_{i-1},x_{i}]$, the probability of $T=t$ is
(12) $\Pr\left(T=t\right)=p_{i}^{t-x_{i-1}}\times(1-p_{i})\times\prod_{j=1}^{i-1}p_{j}^{\Delta_{j}}\,,$
@@ Appendix B Details of the Baseline Methods in Public Dataset Experiments
  • TPM: Tree based Progressive regression Model models video watch time as a series of dependent conditional classification tasks that are arranged into a tree structure. In inference time, the expected video watch time can be obtained by first traversing the tree to find the corresponding leaf node, and then estimating the expectation by a simple model, such as the mean of training samples in that leaf node. Notably, the training objective of TPM explicitly includes the variance of the predicted watch time, as a way to minimize the prediction uncertainty.

### v2

@@ 3.3.2. Simplification: A Stationary Watching Probability
One may simplify the assumption on non-stationary watching probabilities in Section 3.3.1 by assuming a uniform watching probability over the entire video horizon, i.e., the user’s interest in continuing watching the video does not vary with the played time. As will be revealed later on and in Section 3.4, this simplified model is closely related to weighted logistic regression (WLR, Covington et al., 2016).
@@
While Eq. 9 takes the same functional form as the watch-time estimator in WLR (Covington et al., 2016), it operates without extra assumption or approximation in our framework. Concretely, Covington et al. (2016) also estimate watch time by the exponentiated logits (Eq. 9), but require the extra assumption of small click probability to establish certain approximation. Click probability is inapproprate to our target of short video platforms, as discussed in Sec. 1 and 2.1.
@@ 3.4. Connection with WLR
As a final remark, the functional similarity between ${\mathcal{L}}_{\mathrm{WLR}}$ and ${\mathcal{L}}_{\mathrm{v-geo}}$ indicates that WLR approximately models video watch time by a geometric distribution with uniform watching probability over the entire video horizon, which accounts for neither real user’s varying interests in continuing watching nor different video lengths. By contrast, the bucketization strategy employed in our framework can naturally cope with both desiderata should data support. For example, to account for the differences in video length, the model may learn unequal $\{\widehat{p}_{i}\}$ whose values become negligible when the index $i$ is greater than some smaller integers for short videos (due to their smaller lengths) and larger integers for long videos.
@@ 4.1. Public Dataset
CIKM16 Cup dataset contains sequences of anonymous transactions provided by DIGINETICA, and the goal is to predict the dwell time in search results for each session. The dataset contains $310,302$ sessions, a total of $122,991$ items and the average dwell time of each session is $3.981$. For each session, each item in this session is used as a feature in the model input.
@@
We compare our method with state-of-the-art methods for video watch time prediction, including Weighted Logistic Regression (WLR, Covington et al., 2016), Duration-Deconfounded Quantile Regression (D2Q, Zhan et al., 2022), Ordinal Regression (OR, Crammer and Singer, 2001), Tree-based Progressive Regression Model (TPM, Lin et al., 2023), and Classification-Restoration framework with Error Adaptive Discretization (CREAD, Sun et al., 2024). A detailed description of these baseline methods is provided in Appendix B.
@@
where $\mathrm{round}(\cdot)$ denotes rounding to the nearest integer, $c=50$ for the KuaiRec dataset, and $c=100$ for the CIKM dataset, both of which are conveniently chosen based on the scale of the data target values. In inference time, the effect of multiplication by $c$ is reverted before calculating the evaluation metrics. On both datasets, we conveniently use $100$ buckets, defined by the $1$-percentiles of the data target values. Section 4.1.6 conducts ablation study on the number of buckets, showing that our method is generally robust to this design parameter. On the CIKM dataset, both SWaT-Binom and SWaT-Geo use learning rate 2e-3. On the KuaiRec dataset, SWaT-Binom uses learning rate 3e-3 and SWaT-Geo uses 5e-4. These learning rates were selected simply by the converged value of the training loss and the speed of convergence. We use AdamW optimizer (Loshchilov and Hutter, 2017) with weight decay 00.
@@
  • MAE: This metric measures absolute numerical error. Denote the observed target value as $t_{i}$ and the predicted value as $\widehat{t}_{i}$, we have
@@
    where $N$ denotes the test set size. Smaller MAE suggests better model performance.
@@
As a side note, we observed that our models’ performance is stable and roughly the same across multiple random seeds. For better comparison with baseline methods, we followed Lin et al. (2023) and Sun et al. (2024) to not include standard deviations in the result table.
Table 1. Performance of our models on the metrics XAUC and MAE on KuaiRec and CIKM16 datasets. Baseline results are directed cited from Lin et al. (2023) and Sun et al. (2024). The best result in each metric is bold, and the second best result is bold italic.
@@
We compare the results of our models and the baselines in Table 1, where the baseline results are directly cited from Lin et al. (2023) and Sun et al. (2024). In Table 1, our two models SWaT-Binom and SWaT-Geo perform stably across all the metrics and datasets and are almost always ranked top-2. The better performance of our models over the SoTA methods TPM and CREAD indicates the efficacy of our user-centric perspective and statistical framework for modeling video watch time. The gain of our methods over WLR may further testify our consideration on training stability and regularization, which are naturally incorporated in our proposed framework, as discussed in Section 3.4.
@@
From Fig. 4 and Fig. 5 and recall the best baseline results in Table 1, it is clear that our models are generally robust to the number of buckets. While our convenient choice of $100$ buckets in Section 4.1.3 works well across the combination of models, metrics, and datasets, performance on a specific setting may be further improved by tuning the bucket numbers. For example, on the KuaiRec dataset, $50$ buckets suits SWaT-Geo model best, and using $20$ or even $10$ buckets can improve SWaT-Binom model’s XAUC. Using more buckets, such as $200$, however, generally does not provide much gain over using $100$ or fewer buckets. Hence, for a practical rule-of-thumb, one may start with $100$ buckets, and gradually reduce the bucket number while closely monitoring the desired metrics.
(b): Is the training of our models stable and well converged?
@@ 4.2. Industrial Offline Experiment
Due to the large size of the dataset, we train the models with one pass of the training set, i.e., for one epoch. Our models and the baseline use the same model architecture, except for the inevitable changes in the loss function and the corresponding model output layer. All models use the same set of features and hyperparameters, both of which are optimized for the baseline model, which leaves room for further improving our models’ performance. Bucket definition is investigated in Section 4.2.6.
@@
As per industrial common practice, we evaluate the prediction of video watch time by the Pearson correlation between the predicted watch time and the ground truth. Large Pearson correlation indicates that higher predicted values mostly correspond to higher ground-truth values, and thus the prediction model can be more helpful for selecting videos to recommend. As an aside, we re-ran our models several times using data from different time periods, and the performance gains indicated in Table 2 were robust.
@@ 4.2.5. Main Results
Table 2 compares the performance of our three models discussed in Section 3 against the baseline model. It is clear that all our proposals, even the un-bucketized vanilla geometric model, perform significantly better than the baseline, demonstrating the efficacy of our user-centric perspective and statistical framework for modeling video watch time. As a side note, the simplest vanilla geometric model is less performant on other industrial topline metrics, and thus we will continue focusing on our two bucketized models.
@@
We investigate the performance of our SWaT-Binom model and SWaT-Geo model under various bucket definitions. Our intention is to find a bucket definition that performs well while having fewer buckets. We hypothesize that due to the long-tail distribution of video watch time and our desiderata of modeling the right tail well, a dense uniform grid may not be desirable. With this motivation, we compare the following six choices of the bucket endpoints.
@@
Table 3. SWaT-Binom and SWaT-Geo model’s gain over the baseline model, under different choices of bucket endpoints as listed in Section 4.2.6, together with the number of buckets under each choice. The gain is calculated as a direct subtraction by the baseline model’s result in Table 2. Highest numeric in each column is bold.
@@
Table 3 shows the gains of the SWaT-Binom model and SWaT-Geo model over the baseline model under these six definitions of bucket endpoints. We also tabulate the number of buckets under these six definitions. Since bucket endpoints need to be distinct, we remove duplicate values that repeat for different percentiles, and hence the number of buckets can be fewer than the arithmetically calculated counts. For example, in Choice 1, using the 5-percentiles to define the bucket endpoints should ideally give 20 buckets, but removing the duplicates eventually gives us only 18 buckets.
@@ 4.3. Online Experiment
We conduct an online A/B test by implementing our method onto the recommendation system on an real-world short video platform with hundreds of millions of daily-active users. Specifically, our experiment focuses on the performance of our method on long videos, which are more relevant to the event of video watch time.
@@
Table 4. Gain of our model over the production model on long videos in an online A/B test over five days. Reported metrics are accumulated watch time and accumulated watch count. For both metrics, the higher the numeric, the better. The gain is calculated as a direct subtraction by the baseline model’s result.
@@
Table 4 shows the results of online A/B test in each day over five days of traffic, on the metrics accumulated video watch time and accumulated watch counts, both of which are positive metrics, i.e., the higher the better. It is clear that our model provides significant gain on these two metrics in each day of the testing window. This result coincides with our industrial offline experiments (Section 4.2) and public datasets’ experiments (Section 4.1) that our method’s predicted video watch time is more aligned with the ground-truth watch time. Hence, our model can more accurately distinguish videos that attract the users most, and thus the recommendations based on our model increase users’ video watch time and the number of watched videos (video watch count).
@@ 5. Conclusion
Video watch time prediction is one of the central problems in (short) video recommendation, as it directly affects the recommendation quality and user experience. Unlike conventional binary prediction tasks, this problem is significantly challenged by the complex user-video interaction; while most of the existing literature focus on black-box mechanical enhancement of the classical regression or classification losses. In this paper, we for the first time take on a user-centric perspective to model video watch time, from which we propose a white-box statistical framework that translates our domain knowledge on users’ video-watching behavior into principled statistical models. To better model users’ fluctuating interests in continuing watching the video over the video horizon, we employ the bucketization strategy, which additionally helps to stabilize training objectives. Extensive experiments of various nature validate our perspective and derived models, showing that our predictions align well with the ground-truth video watch time.
6. Limitation and Future Work
As an earliest effort in modeling video watch time through user behavior analysis, our paper biases towards succinct statistical models derived from relatively strict assumptions, such as Assumption 1 in Section 3.2, so as to train and deploy models efficiently. Nevertheless, our competitive experiment results validate our unique direction of modeling video watch time from the user-centric perspective and open the door for driving new statistical watch time models by softening and sophisticating those assumptions.
With accurate video watch time prediction, our SWaT models can be directly connected with explicit duration bias mitigation. For example, one may utilize SWaT model’s predictions without being overly sensitive to large predicted values, such as making video recommendations based on $\log(\text{predicted watch time}+1)$ rather than the raw predicted watch time; or use our SWaT framework to estimate a baseline model to contrast for duration debiasing.
Notice that our SWaT-Binom and SWaT-Geo are derived by respective user behavior assumptions, and that industrial recommendation platform typically contains various types of users. One may therefore cluster users and apply SWaT-Binom and SWaT-Geo correspondingly to each user segment, and/or use (learnable) personalized bucketization for each user, potentially with the help of reinforcement learning (Yang et al., 2022b, a, c) or reinforcement learning from human/AI feedback (Feng et al., 2023; Yang et al., 2023, 2024).
Acknowledgements.
We’d like to thank Ali Selman Aydin, Hongjie Bai, Vladimir Batygin, Jack Chai, Assaf Cohen, Shilin Ding, Yuhao Du, Bo Feng, Gavin Feuer, Chenzhang He, Wenjie Hu, Xinyao Hu, Cheng Huang, Gong Ke, Romaine Knight, Chunlei Li, Rui Li, Sophia Liang, Yijia Liu, Matt Ma, Penny Pan, Jing Qian, Yisong Song, Chris Steger, Dayong Wang, Meihong Wang, Zellux Wang, Zhe Wang, Chuanqi Wei, Yanhong Wu, Hong Yan, Lei Yuan, Hui Zhang, Lizhu Zhang, Mengchee Zhang, Qunshu Zhang, and Yuting Zhang for their general support of the project.
@@ Appendix A The Expectation of SWaT-Geo Model
Recall from Section 3.3.1 that for a given set of continuation probabilities $\left\{p_{i}\right\}_{i=1}^{N+1}$, if the observation $t\in B_{n}=(x_{n-1},x_{n}]$, the probability of $T=t$ is
(12) $\Pr\left(T=t\right)=p_{n}^{t-x_{n-1}}\times(1-p_{n})\times\prod_{i=1}^{n-1}p_{i}^{\Delta_{i}}\,,$
@@ Appendix B Details of the Baseline Methods in Public Dataset Experiments
  • TPM: Tree-based Progressive regression Model models video watch time as a series of dependent conditional classification tasks that are arranged into a tree structure. In inference time, the expected video watch time can be obtained by first traversing the tree to find the corresponding leaf node, and then estimating the expectation by a simple model, such as the mean of training samples in that leaf node. Notably, the training objective of TPM explicitly includes the variance of the predicted video watch time, as a way to minimize the prediction uncertainty.